OpenStudy (anonymous):

establish the reductions formula using the sum and difference identities. 1. sin((3pi/2)- x)= -cos x 2. cot((-7pi/2) - x)= tan x

hartnn (hartnn):

u know the sum and diff. identities ?

OpenStudy (anonymous):

nope

hartnn (hartnn):

ok, sin (A-B) = sin A cos B -cos A sin B now put, A=3pi/2, B=x so, sin(3pi/2-x)=sin (3pi/2) cos x - cos (3pi/2) sin x =.................... can u solve further ?

hartnn (hartnn):

know the values of sin(3pi/2) and cos (3pi/2) ?

OpenStudy (anonymous):

what is the value?

hartnn (hartnn):

sin(3pi/2)=-1, cos(3pi/2)=0. put these values and tell me what u get as right side......

OpenStudy (anonymous):

should the answers the exact value?

hartnn (hartnn):